OpenStudy (anonymous):

Nathaniel can weld a railing in 75 minutes. Brenda can weld a railing 25 minutes faster. If they work together, how many minutes does it take them to weld the railing?

OpenStudy (amoodarya):

\[N=\frac{1}{75} per \space \min\\B=\frac{1}{75-25}per \space \min\\N+B =\frac{1}{75}+\frac{1}{50}\\final \space answer =\frac{1}{\frac{1}{75}+\frac{1}{50}}\]

OpenStudy (anonymous):

So what is the answer if the choices are 25 minutes 30 minutes 50 minutes 150 minutes

OpenStudy (amoodarya):

1/75+1/50=1/30 so ...?

OpenStudy (anonymous):

150 is the cd...sooo hold on

OpenStudy (anonymous):

I got 5/5....

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

50

OpenStudy (amoodarya):

\[\frac{1}{75}+\frac{1}{50}=\frac{1}{25}(\frac{1}{3}+\frac{1}{2})=\frac{1}{30}\\\frac{1}{\frac{1}{75}+\frac{1}{50}}=\frac{1}{\frac{1}{30}}=30\]

OpenStudy (anonymous):

Am I right?

OpenStudy (amoodarya):

yes